OpenStudy (anonymous):

Solve the equation explicitly: dx/dt - x^3 = x

OpenStudy (anonymous):

I had to cheat to recall how to do some of this. The first part is to change this into a workable integral. We need the dt by itself: \[\frac{dx}{dt}-x^3=x\] \[\frac{dx}{dt}=x+x^3\] \[\frac{dx}{x+x^3}=dt\] Take the integral of both sides: \[\int\limits dt= \int\limits (x+x^3)\]

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integral+1%2F%28x%2Bx%5E3%29 See the show steps. I'm a little bit to rsuty to recall the specifics to type them out.

OpenStudy (anonymous):

And that integral should be: \[\int dt = \int (x+x^3)dx\]